Simons Center for Geometry and Physics Arts Program

Report to Board of Trustees

May 21, 2013

1. Administration. Since January 2012 the SCGP Arts Program has been without a Director. Exhibitions have been curated by free-lance personnel, coordinated by the Advisory Committee and administered by Elyce Winters.

The exhibits, concerts and events have been successful. Nevertheless, the lack of direction has meant that the SCGP Arts Program is not realizing its potential, especially when it comes to

• integrating our activity with that of the various Stony Brook departments and students,

• communicating, to the entire campus population and to the surrounding community, what the Center has to offer.

At the last Trustees meeting, Provost Assanis suggested that the (half-time) position of SCGP Arts Program Director/Curator could be combined with a half-time position in one of the SBU departments. On May 18 we heard from John Lutterbie, Chair of Art/Theatre Arts, that he and the Provost had discussed this option and tentatively agreed to proceed with the recruitment of a 5-year Visiting Scholar, 50% in SCGP and 50% in Art. In my opinion it would be excellent for us if this search could start as soon as possible.

Even with a search starting immediately, the Arts Program cannot count on having a Director before January or maybe September 2014. I would like to recommend that a oneyear half-time appointment be made as interim program director. There are several very good candidates available. This would allow programming for the 2013-2014 academic year to start right away.

In this connection, I hope the Trustees will consider increasing the SCGP Arts Program budget to cover this temporary appointment and our half of the Visiting Scholar, when that position is filled. The $100,000 currently allocated turns out to be just sufficient for the cost of mounting four or five exhibitions a year (including transportation costs, insurance, curatorial fees, receptions, overnight or 2-day lodging for artists and invited speakers) along with a modest program of concerts during the year and the summer workshop. Half of the salary for a professional at the level we should be seeking will be in the $35,000-$40,000 range, in my estimation.

2. Artist-in-Residence Program. I believe that the Artist-in-residence component of the SCGP Arts Program has not worked out as planned. The ideal of an artist collaborating with Center scientists to create works of art of high aesthetic and scientific quality now seems to have been unrealistic. On the other hand, workshops organized in conjunction with an exhibition (for example origami, 3D printing) have been very successful. I propose that our Artist-in-Residence be terminated after this summer.

3. The current SCGP Arts Program Advisory Committee includes members of the Stony Brook campus and community who have experience with the interaction between art, science and technology:

Stephanie Dinkins, Art Department

George Hart, computer scientist and sculptor

John Lutterbie, Theatre Arts and Art Departments

Tony Phillips, Mathematics Department, Chair

Daniel Weymouth, Music Department

Elyce Winters and I plan to review this committee over the summer: involve new people, more people and more from other departments (including Engineering and non-Fine-Arts Humanities), and at least two community non-academics with experience in the arts and curation (we are thinking of Philip Palmedo and Judith Levy).

4. Please try to visit the current exhibition: Paths: Charting, Navigating & Bridging, curated by Alastair Noble. It is an excellent example of the kind of show we can put on here: reasonably low-budget, visually engaging and with a strong scientific (geometric in this case) organizing principle.

Respectfully submitted,

Tony Phillips, Arts Program Advisory Committee Chair